OpenStudy (enricoanders):

Two cyclists, 96 miles apart, start riding toward each other at the same time. One cycles 2 times as fast as the other. If they meet 4 hours later, what is the speed (in mi/h) of the faster cyclist? a. Write an equation using the information as it is given above that can be solved to answer this problem. Use the variable r to represent the speed of the slower cyclist. b. What is the speed of the faster cyclist?

OpenStudy (enricoanders):

@satellite73

OpenStudy (danjs):

One travels speed r, the other is 2*r, overall the distance is being covered at r+2r = 3r, the sum of their speeds.

OpenStudy (danjs):

Distance = rate*time 96 = 3*v*4

OpenStudy (enricoanders):

so thats the equation

OpenStudy (danjs):

yes, d = ( r + 2r)*t 96 = (r + 2r) * 4 96 = 12r r=8 mi/h

OpenStudy (enricoanders):

okay thanks